OpenStudy (anonymous):

Solve the following system of equations by using the elimination method. 2x - 3y = -10 3x + 5y = 4

OpenStudy (anonymous):

Multiply 1st line by -3 and second line by 2. This causes the x's to cancel each other out.

OpenStudy (anonymous):

Therefore you can solve for y, once you know y, u can solve x using any of the two equations given.

OpenStudy (anonymous):

@ironictoaster can you do a step by step i get confused once i finish multiplying by -3 and 2

OpenStudy (anonymous):

Multiplying the 1st line out by -3 2x(-3)-3y(-3)=-10(-3) -6x + 9y = 30 Multiplying the 2nd line out by by 2 3x(2) + 5y(2) = 4(2) 6x + 10y = 8

OpenStudy (anonymous):

New simultaneous equations are below. Notice if you add the x's together it gives zero. We add can add the y's together and solve for y. Once you got y, you sub in the answer into any of the two equations and solve for x -6x + 9y = 30 6x + 10y = 8

OpenStudy (anonymous):

So you'll have 19y = 38 and y=2

OpenStudy (anonymous):

Sub in 2 for y in any equation and solve for x!

OpenStudy (anonymous):

@ironictoaster ok thankyou so much dude!!!! great explanation!!!!!